### v1

# Pair-breaking as the fundamental limit to persistent-current stabilization in fermionic superfluids

Abstract
We study the stability of persistent currents in fermionic superfluids with impurities within the BCS regime by using time-dependent density functional theory. Unlike in Bose–Einstein condensates, we find that current stabilization by impurities is intrinsically limited by the pair-breaking threshold. Below the threshold, impurities enhance winding number stability, but pair-breaking continues to drive dissipation of the flow. Above this critical velocity, superflow destabilizes, emitting vortices. Impurities then govern vortex mobility and pinning, exhibiting regimes of collective pinning and hopping. Moreover, pinned vortices do not guarantee dissipationless flow due to ongoing pair-breaking. Our results identify pair breaking as the fundamental mechanism that sets the ultimate limit of persistent-current stability in fermionic superfluids, providing insights relevant to ultracold Fermi gases and neutron-star matter.
Introduction. — The flow of a Fermi superfluid through a spatially inhomogeneous medium plays a central role in many physical systems and poses several unresolved theoretical challenges. In such environments, the superfluid fraction is generally reduced below unity, implying that only part of the fluid participates in coherent, collective motion Leggett (1970); Orso and Stringari (2024); Saslow (1976); Galli et al. (2007); Biroli et al. (2011); Pérez-Cruz et al. (2025). The remaining fraction, effectively decoupled from the superflow, is expected to behave as a normal component even at zero temperature, although it may not contribute to entropy production. Whether this decoupling occurs depends sensitively on both the nature of the inhomogeneities and the velocity of the superfluid. A systematic classification of the dynamical regimes arising under these conditions remains an open problem.
Impurities play a crucial role in superfluid dynamics. Much like in classical fluids, they can give rise to dissipative processes. In Bose superfluids, impurities often serve as nucleation sites for quantum vortices Sasaki et al. (2010); Christenhusz et al. (2025), which are the primary agents coupling the superfluid and normal components and enabling energy transfer between them. For instance, the Hall-Vinen-Bekarevich-Khalatnikov model (a generalization of Landau’s two-fluid theory) requires both a superfluid fraction below unity and the presence of quantized vorticity for dissipation to occur during the dynamics (see Barenghi et al. (2006) for reviews). Thus, impurities influence superfluid behavior in two fundamental ways: first, their mere presence effectively reduces the superfluid fraction; second, by seeding vortices, they enable the transfer of flow energy into the normal component, which subsequently dissipates it into internal excitations.
Ultracold atomic gases provide a tunable platform for studying superfluid transport. In two-dimensional annular or toroidal traps, persistent currents arise with quantized circulation determined by the winding number $w$ Ryu et al. (2007); Cai et al. (2022); Pecci et al. (2021); Del Pace et al. (2022); Polo et al. (2025). In Bose–Einstein condensates (BECs), impurities or barriers can nucleate quantized vortices Ramanathan et al. (2011); Williams et al. (2010); Wright et al. (2013); Kwon et al. (2015); Gan et al. (2025); Pezzè et al. (2024); Xhani et al. (2023); Del Pace et al. (2022); Moulder et al. (2012); Kumar et al. (2017), leading to phase slips and the decay of superflow. Interestingly, recent studies have shown that introducing a controlled number of weak links (Josephson junctions) can, paradoxically, stabilize the superflow by increasing the critical current required for decay Pezzè et al. (2024). This effect arises due to the reduction of phase jumps or maximum velocity at the weak-links with increasing junction number. Hence, engineered weak links can suppress phase slips and enhance the metastability of persistent currents.
An open question is whether this behavior is a universal feature of all superfluids or specific to BECs. Fermi superfluids constitute a distinct class of systems characterized by additional complexity. Unlike BECs, superfluidity in these systems arises from the formation of Cooper pairs, which become fragile as one approaches the so-called BCS regime: a regime of weak interactions. Pair-breaking processes reduce the superfluid fraction and can trigger vortex nucleation, leading to dynamics that differ markedly from those of BECs Xhani et al. (2025a); Pisani et al. (2024); Chen et al. (2025). In this work, we show that pair breaking imposes a fundamental limit on the stability of persistent currents: beyond a critical velocity associated with the pair breaking, stability cannot be enhanced by introducing impurities, unlike in the case of BECs Xhani et al. (2025b); Pezzè et al. (2024). This mechanism is analogous to the depairing current known from superconductivity Kunchur (2019); MacManus-Driscoll et al. (2004); Miura et al. (2022, 2024), which sets the intrinsic upper bound on supercurrent achievable in superconductors. Here, we identify its counterpart for superfluids and analyze how stability depends on impurity size and spacing, as well as on their influence on vortex mobility in Fermi superfluid rings within the BCS regime.
Theoretical model. — We study the dynamics of persistent currents in a weakly attractive fermionic superfluid within the Bardeen–Cooper–Schrieffer (BCS) regime. The evolution is described using time-dependent density functional theory for superfluid Fermi gases Bulgac (2007); Bulgac and Forbes (2008); Bulgac et al. (2012), in the form of the extended Superfluid Local Density Approximation (SLDAE) Boulet et al. (2022). This framework has been successfully applied to Josephson dynamics Wlazłowski et al. (2023), persistent currents Xhani et al. (2025a), vortex dynamics Barresi et al. (2023); Bulgac et al. (2011); Wlazłowski et al. (2018), Higgs modes Bulgac and Yoon (2009), spin-imbalanced systems Bulgac et al. (2006); Bulgac and Forbes (2008); Kopyciński et al. (2021), and quantum turbulence Wlazłowski et al. (2015); Hossain et al. (2022).
The time-dependent SLDAE equations are formally equivalent to the time-dependent Bogoliubov–de Gennes (BdG) equations (with $\hbar=m=1$):
$\begin{pmatrix}h(\vec{r},t)&\Delta(\vec{r},t)\\ \Delta^{*}(\vec{r},t)&-h^{*}(\vec{r},t)\end{pmatrix}\begin{pmatrix}u_{n}(\vec{r},t)\\ v_{n}(\vec{r},t)\end{pmatrix}=i\frac{\partial}{\partial t}\begin{pmatrix}u_{n}(\vec{r},t)\\ v_{n}(\vec{r},t)\end{pmatrix}$, (1)
where $(u_{n},v_{n})^{T}$ are the Bogoliubov amplitudes. Unlike the conventional BdG formalism valid only for weak coupling, SLDAE accurately captures the entire BCS–BEC crossover. Here we focus on $a_{s}k_{F}=-1$, with $k_{F}=(3\pi^{2}\rho_{0})^{1/3}$ defined by the bulk density $\rho_{0}$, and $a_{s}$ being the scattering length. The normal, anomalous, and current densities are expressed as
$\rho=2\sum_{n}\lvert v_{n}\rvert^{2},\,\,\,\nu=\sum_{n}u_{n}v_{n}^{*},\,\,\,\vec{j}=2\sum_{n}\textrm{Im}\left(v_{n}\vec{\nabla}v_{n}^{*}\right)$ (2)
where the sums run over quasiparticle states with $E_{n}\in(0,E_{c})$, with the cutoff $E_{c}$ that regularizes ultraviolet divergences Bulgac and Yu (2002); Bulgac (2007). The single-particle Hamiltonian and pairing field are
$h=-\tfrac{1}{2}\nabla^{2}+U+V_{\textrm{ext}}(\vec{r},t),\qquad\Delta=-\frac{C}{\rho^{1/3}}\nu,$ (3)
with $U$ the mean-field potential, $V_{\textrm{ext}}$ the external potential (trap and impurity potentials), and $C$ a regularized pairing constant. Their explicit expressions are provided in Ref. Boulet et al. (2022) and assure proper reproduction of the quantum Monte Carlo equation of state and pairing gap. For a uniform system at a considered coupling constant, SLDAE gives $\Delta/\varepsilon_{F}\approx 0.134$, where $\varepsilon_{F}=k_{F}^{2}/2$ is the Fermi energy, and the corresponding coherence length is $\xi=\varepsilon_{F}/k_{F}\Delta\approx 7.5\,k_{F}^{-1}$.
The superfluid is confined in a two-dimensional ring potential with inner and outer radii $R_{\mathrm{in}}=30k_{F}^{-1}$ and $R_{\mathrm{out}}=60k_{F}^{-1}$, respectively. Trapping is applied only in the xxyy plane, while translational symmetry is imposed along the $z$ direction; thus, the system is effectively three-dimensional at the microscopic level. A controllable superflow is generated using the phase-imprinting technique. Specifically, a phase $\phi=w_{0}\arctan(x,y)$ is imprinted onto the order parameter, $\Delta(x,y)=|\Delta(x,y)|e^{i\phi(x,y)}$, and the corresponding ground state is obtained from stationary simulations $i\partial/\partial t\rightarrow E_{n}$ in Eq. (1)] in the absence of impurities. This procedure initializes a rotating superfluid with a total phase winding of $2\pi w_{0}$, corresponding to a persistent-current state with winding number $w_{0}$. A key distinction between BEC and BCS superfluids arises already at this stage: in the latter, the winding number cannot exceed the pair-breaking threshold $w_{\mathrm{pb}}\approx 2R_{\mathrm{in}}v_{\textrm{pb}}$, where $v_{\textrm{pb}}=\sqrt{\sqrt{\mu^{2}+\Delta^{2}}-\mu}$ and $\mu$ is the chemical potential Xhani et al. (2025a). Above this limit, pair breaking suppresses the order parameter near the inner radius, which may lead to vortex emission and subsequent decay of the current. For the present setup, the estimated threshold is $w_{\mathrm{pb}}=5$.
Figure 1: (a) Magnitude of the pairing field $|\Delta(x,y)|$ for $w_{0}=7$ in the absence of impurities ($N_{d}=0$) and with $N_{d}=6$ impurities, shown at $t\varepsilon_{F}\approx 400$. The impurity size is $s_{2}=0.67\xi$. The imposed flow direction and vortex positions are indicated by the swirling arrows. (b) Flow–energy decay rate $A/\tau$ for the impurity size $s_{1}=1.33\xi$ as a function of the winding number $w_{0}$ and the number of impurities $N_{d}$. The grey markers separate the unstable regime, where the winding number decays, from the stable regime, where $w(t)$ remains constant. The dashed black line marks the pair-breaking threshold $w_{\mathrm{pb}}$ for this configuration.
Similar to the experimental procedure Pezzè et al. (2024), impurities are introduced dynamically during the simulation. Each impurity is modeled by a Gaussian potential symmetrically distributed along a ring (forming a circular lattice):
$V_{\textrm{imp}}(\vec{r},t)$ $=V_{0}(t)\sum_{i=1}^{N_{d}}\exp\!\left[\frac{(x-x_{i})^{2}+(y-y_{i})^{2}}{2\sigma^{2}}\right],$ (4)
where $V_{0}(t)$ is the time-dependent amplitude, ramped from zero to its final value ($\approx 2\varepsilon_{F}$) within the time interval $(0\text{–}25)\varepsilon_{F}^{-1}$. The impurity centers are positioned at $x_{i}=R_{0}\cos(2\pi i/N_{d})$ and $y_{i}=R_{0}\sin(2\pi i/N_{d})$, with $R_{0}=(R_{\mathrm{in}}+R_{\mathrm{out}})/2$. We explore the system’s response as a function of two control parameters: the number of impurities $N_{d}$ and their size $s$, defined as the full width at half maximum, $s\approx 2.355\sigma$. Two representative regimes are considered: impurities larger than the coherence length ($s_{1}=1.33\xi$) and smaller impurities ($s_{2}=0.67\xi$). In our simulations, we vary $2\leq N_{d}\leq 16$.
The equations of motion (1) are solved numerically on a computational grid of size $128\times 128$, with lattice spacing $dxk_{F}=1$, and plane waves along the third direction are assumed. All simulations are performed using the W-SLDA Toolkit WSL .
Results — Representative snapshots from the time-dependent simulations are shown in Fig. 1(a) for the case $w_{0}=7$, i.e., above the pair-breaking threshold, and for two configurations: without impurities ($N_{d}=0$) and with $N_{d}=6$ impurities. As expected, for the case of no impurities, the persistent current is unstable. Vortices are spontaneously nucleated (at the inner edge for $N_{d}=0$), which leads to the decay of the current. Systematic simulations confirm that, for $w_{0}>w_{\mathrm{pb}}$, the persistent current cannot be stabilized regardless of the number of impurities introduced. This behavior contrasts sharply with observations Xhani et al. (2025b); Pezzè et al. (2024) in the BEC regime, where the addition of impurities or Josephson junctions can enhance the stability of the persistent current. This fundamental difference between bosonic and fermionic superfluids constitutes one of the main conclusions of this work.
To make this statement quantitative, we introduce an energy-based diagnostic by computing the flow energy, which measures the total current in the system:
$E_{\mathrm{flow}}(t)=\int\frac{\vec{j}^{2}(\vec{r},t)}{2\rho(\vec{r},t)}\,d^{3}\vec{r}.$ (5)
This quantity includes contributions from both the superfluid and normal components, since within the DFT framework $\vec{j}\neq\rho\,(\hbar/2)\nabla\phi$ (the factor $2$ corresponds to the Cooper pair mass). This differs from, for example, the Gross–Pitaevskii approach, where the equality holds exactly. Consequently, $E_{\mathrm{flow}}$ captures both vortex dynamics and normal-fluid dissipation Xhani et al. (2025a). We also monitor the condensation energy,
$E_{\mathrm{cond}}(t)=\frac{3}{8}\int\frac{|\Delta(\vec{r},t)|^{2}}{\varepsilon_{F}(\vec{r},t)}\rho(\vec{r},t)\,d^{3}\vec{r},$ (6)
where $\varepsilon_{F}(\vec{r})$ is the local Fermi energy Wlazłowski et al. (2023). This quantity measures the pairing strength and provides a direct probe of pair-breaking processes during the evolution. To quantify dissipation, the temporal profile of the flow energy, $E_{\mathrm{flow}}(t)$, is fitted with an exponential function, $f(t)=A\exp\left[-(t-t_{0})/\tau\right]+B$ where $A$ and $B$ are related to initial and final values, $\tau$ the decay time, and $t_{0}=25\,\varepsilon_{F}^{-1}$ denotes the time when the impurities are fully ramped on. To characterize the topological stability of the current, we also define the critical winding number $w_{c}$ as the largest initial winding $w_{0}$ that remains stable over the time interval, precisely for $t\varepsilon_{F}\leq 1000$.
Fig. 1(b) summarizes the dependence of $w_{c}$ on the number of impurities $N_{d}$, for impurity size $s_{1}=1.33\xi$. The critical winding increases with $N_{d}$, consistent with observations for BECs Xhani et al. (2025b); Pezzè et al. (2024), until it saturates at the pair-breaking threshold $w_{\mathrm{pb}}=5$. For the smaller defect size ($s_{2}=0.67\xi$), we observe that $w_{c}=w_{\mathrm{pb}}$ holds throughout the entire range of considered $N_{d}$. Thus, $w_{\mathrm{pb}}$ sets an upper bound on the current’s stability. We note that this threshold depends on the inner ring radius and could be increased by choosing a larger $R_{\mathrm{in}}$. To illustrate this behavior, we focus, for example, on the case $w_{0}=3$ as a function of $N_{d}$. As the number of impurities increases, the flow evolves from an unstable regime, where the winding decays for $N_{d}\leq 2$, to a stable regime with constant $w(t)$ for $N_{d}>2$. However, stabilization of the winding number does not necessarily imply dissipationless flow. This is evident from the color map, which shows the decay rate $A/\tau$: values near zero correspond to nearly dissipationless flow, while finite $A/\tau$ emerge even when $w(t)$ stabilizes. Hence, topological stability ($w_{c}$) and energetic stability of the current are distinct phenomena.
Figure 2: Dissipation mechanisms of persistent currents in the presence of impurities. (a) Flow–energy decay rate $A/\tau$ obtained from exponential fits for impurity sizes $s=1.33\xi$ (main panel) and $s=0.67\xi$ (inset). (b) Time evolution of the winding number $w(t)$ for $s=1.33\xi$. (c) Per-particle flow–energy change $\delta E_{\mathrm{flow}}/N$ between initial and final states. (d) Condensation–energy change $\delta E_{\mathrm{cond}}$ highlighting pair-breaking effects. Panels (c) and (d) show that the impurity size primarily governs variations in the flow and condensation energies, rather than the topological properties of the order parameter, which in turn are encoded in $w(t)$.
To further quantify dissipation, we evaluate the change in the flow energy per particle, $\delta E_{\mathrm{flow}}/N=[E_{\mathrm{flow}}(t_{\textrm{f}})-E_{\mathrm{flow}}(t_{\textrm{i}})]/N\varepsilon_{F}$ and the relative change in the condensation energy $\delta E_{\mathrm{cond}}=[E_{\mathrm{cond}}(t_{\textrm{f}})-E_{\mathrm{cond}}(t_{\textrm{i}})]/E_{\mathrm{cond}}(t_{\textrm{i}})$, where $t_{\textrm{i}}=25\varepsilon_{F}^{-1}$ and $t_{\textrm{f}}=1000\varepsilon_{F}^{-1}$. Fig. 2 presents these quantities for $w_{0}=3$ and $w_{0}=7$ (below and above $w_{\mathrm{pb}}$) and for both impurity sizes. The impurity size plays a crucial role: smaller impurities produce substantially lower losses in both $E_{\mathrm{flow}}$ and $E_{\mathrm{cond}}$. For subcritical windings ($w_{0}=3<w_{\mathrm{pb}}$), increasing the impurity number $N_{d}$ enhances dissipation and accelerates decay. For the range $2<N_{d}\leq 16$, the winding $w_{0}=3$ remains below or equal to $w_{c}$, so no vortices are emitted, $w(t)$ stays constant, and dissipation arises solely from pair breaking, which intensifies with impurity density irrespective of impurity size. When $w_{0}>w_{c}$, vortices appear and their number increases with both $w_{0}$ and $N_{d}$. At these higher windings, dissipation originates from the combined effects of vortex emission and pair breaking. For instance, at $w_{0}=7$ (exceeding both $w_{\mathrm{pb}}$ and $w_{c}$), vortices are continuously nucleated, independent of $N_{d}$ or impurity size $s$, which now constitutes the dominant dissipative channel.
Figure 3: (a,b) Radial and angular vortex mobilities as functions of the number of impurities $N_{d}$ for two impurity sizes, at $w_{0}=7$. (c) Two-dimensional vortex trajectories for $w_{0}=7$ for different numbers of impurities and their sizes.
Next, we investigate the relationship between vortex mobility and dissipation by analyzing vortex dynamics and the impact of impurities on vortex mobility. The radial and angular mobilities, shown in Fig. 3, are defined as
$v_{r}$ $=\frac{\overline{r}}{N_{v}}\sum_{k=1}^{N_{v}}\frac{\Delta r_{k}}{\Delta t_{k}},\quad v_{\theta}=\frac{\overline{r}^{2}}{N_{v}}\sum_{k=1}^{N_{v}}\frac{\Delta\theta_{k}}{\Delta t_{k}},$ (7)
where $N_{v}$ is the number of vortices, $\overline{r}=\frac{1}{N_{v}}\sum_{k}r_{k}^{(\textrm{i})}$ is the mean initial radius, and $\Delta r_{k}=r_{k}^{(\textrm{f})}-r_{k}^{(\textrm{i})}$, $\Delta\theta_{k}=\theta_{k}^{(\textrm{f})}-\theta_{k}^{(\textrm{i})}$, $\Delta t_{k}=t_{k}^{(\textrm{f})}-t_{k}^{(\textrm{i})}$ denote the radial, angular, and temporal displacements between the initial (i) moment of vortex nucleation and its annihilation or the end of the simulation (f). These results correspond to a fixed initial winding number $w_{0}=7$. Impurities significantly impact vortex emission and motion, with trajectories influenced by the size and spacing of the impurities. In the absence of impurities, vortices are symmetrically emitted along the inner ring, forming a periodic pattern that propagates freely. Adding impurities modifies the number and dynamics of vortices. Weak impurities usually deflect vortex paths, while stronger ones induce pinning and trigger the release of additional vortices, visible as color variations in Fig. 3c) for $s_{1}=1.33\xi$. As impurity density increases, the radial mobility exhibits contrasting trends for the two impurity sizes. For larger impurities, both radial and angular mobilities decrease steadily with $N_{d}$, reflecting stronger collective pinning (see $N_{d}=6$ in Fig. 3c). For smaller impurities, $v_{r}$ initially grows (up to $N_{d}=6$), and then starts to decrease. Contrary to conventional expectations, these results show that even when vortices are pinned to impurities (immobile), the current or flow energy remains dissipative, see Fig. 2c). This dissipation arises from pair-breaking processes induced by the impurities. Notably, we find a signature of stabilization effect on the flow energy, for small defects and moderate impurity concentrations ($N_{d}>6$), visible as a plateau in $\delta E_{\mathrm{flow}}$ for $s_{2}=0.67\xi$ and $w_{0}=7$. Interestingly, at $N_{d}=16$ and larger impurity size, vortices exhibit tunneling between impurity sites ($v_{\theta}$ saturates at a non-zero value), signaling a distinct dynamical regime.
Conclusions — We have investigated the stability of persistent currents in fermionic superfluids within the BCS regime using time-dependent density functional theory. Our results demonstrate that the stability of supercurrents is fundamentally limited by a pair-breaking threshold, in sharp contrast to Bose–Einstein condensates, where impurity-induced stabilization is possible Xhani et al. (2025b); Pezzè et al. (2024).
Two distinct channels of energy dissipation emerge in fermionic systems: (i) topological relaxation, associated with winding-number decay and vortex proliferation above the critical winding $w_{c}$, and (ii) microscopic pair-breaking processes, which operate even below this threshold. The latter channel is primarily governed by impurity size, as larger impurities enhance the local suppression of the order parameter and accelerate energy loss without necessarily inducing vortex nucleation. In the vortex-nucleation regime, the size and spacing of impurities strongly influence vortex mobility. Larger impurities reduce vortex mobility, indicating a transition to a pinning regime, while closely spaced impurities allow for vortex hopping between pinning sites. These findings provide a microscopic picture of how disorder and geometry jointly control the interplay between dissipation and stability in fermionic superflows.
The results are directly relevant to astrophysical systems such as neutron stars. Although neutron matter is strongly interacting, its pairing gap and coherence length correspond to a weak-coupling BCS regime Carlson et al. (2012). The impurity spacings explored here are comparable to the expected distances between nuclear clusters in the neutron-star crust Negele and Vautherin (1973); Chamel (2012). Hence, the mechanisms identified (pair-breaking dissipation and impurity-mediated pinning) offer insight into vortex dynamics in neutron-star superfluids and their role in phenomena such as pulsar glitches Chamel and Haensel (2008); Haskell and Melatos (2015); Antonopoulou et al. (2022).
Note added in proof. A recent work Xhani et al. (2025b) has shown increased current stability with impurity density and its dependence on the impurity arrangement with bosonic superfluids.
I Acknowledgements
We thank Brynmor Haskell for fruitful discussions. This work was financially supported by the (Polish) National Science Center Grants No. 2022/45/B/ST2/00358. K. X. acknowledges funding from the Italian MUR (PRIN DiQut Grant No. 2022523NA7). We acknowledge the Polish high-performance computing infrastructure, PLGrid, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland), and the LUMI consortium through PLL/2024/07/017603.
Author contributions:
The work was conceived by K.X. Calculations were executed by A.B., and data analysis was performed by B.T. All authors contributed to research planning, interpretation of the results, and manuscript writing.
Data avalibility
The data that support the findings of this article are openly available Tüzemen et al. (2025).